The base of a ladder is 10 feet fro the wall. The top of the ladder is 11 feet from the floor. Find the length of the ladder to the nearest thousandth.
Answer by ankor@dixie-net.com(22740)

The base of a ladder is 10 feet fro the wall. The top of the ladder is 11 feet from the floor. Find the length of the ladder to the nearest thousandth.
:
A typical pythagorus problem where a^2 + b^2 = c^2
:
In this problem a = 10, b = 11, c = length of the ladder
:
c^2 = 10^2 + 11^2
c^2 = 100 + 121
c^2 = 221
c = sqrt(221)
c = 14.866 ft is the length of the ladder